OpenStudy (anonymous):

in a two-candidate election 1401 votes were cast. if 30 voters had switched their votes from the winner to the loser, the loser would have won by 5 votes. How many votes did each candidate actually receive?

OpenStudy (lgbasallote):

let x be the votes won by candidate A y be the votes won by candidate B total votes is 1401 so x + y = 1401 if 30 votes switched from winner to loser then the winner would have won by 5 votes let's say x won and y lost so x - 30 = y+ 35 because x lost 30 votes and they go to y then why wins by 5 votes. i can rewrite this as x - y = -5 if i subtract y from both sides and add 30 to both sides now let's do the system x + y = 1401 x - y = -5 add them 2x = 1396 divide both sides by 2 x = 698 so y has 703 this would be the one with the edited votes...to get the original just add 30 to x and subtract 30 from y so x = 728 y = 673

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

is there a way to write that into 1 equation